# Fermi-liquid transport beyond the upper critical field in superconducting La2PrNi2O7 thin films

### v1

@@ Acknowledgements
We thank Toni Helm for experimental support with pulsed-field measurements and Bai Yang Wang for stimulating discussions. We also thank Mark Gibson and Jiarui Li for assistance in sample preparations. This work is supported by the Yushan Fellow Program (MOE-112-YSFMS-0002-002-P1) and the Center for Quantum Science and Technology (CQST) within the framework of the Higher Education Sprout Project by the Ministry of Education (MOE), Taiwan, and by the National Science and Technology Council, Taiwan (NSTC 113-2112-M-008-044-MY3) (Y.-T.H). This work is also supported by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division under contract no. DE-AC02-76SF00515 (synthesis and transport measurements) as well as SuperC and the Kavli Foundation (in situ transport) (Y.L., Y.T., Y.Y., H.Y.H.), and by Grants-in-Aid for Scientific Research (KAKENHI, grant numbers 22H00104 and 25H00600) from the Japanese Society for the Promotion of Science (JSPS) (Y.K., V.S.). We further acknowledge support of the HLD at HZDR, member of the European Magnetic Field Laboratory (EMFL) as well as support under the European Union’s Horizon 2020 research and innovation programme through the ISABEL project (No. 871106).

### v2

@@ Acknowledgements
We thank Toni Helm for experimental support with pulsed-field measurements. We also thank Mark Gibson and Jiarui Li for assistance in sample preparations. This work is supported by the Yushan Fellow Program (MOE-112-YSFMS-0002-002-P1) and the Center for Quantum Science and Technology (CQST) within the framework of the Higher Education Sprout Project by the Ministry of Education (MOE), Taiwan, and by the National Science and Technology Council, Taiwan (NSTC 113-2112-M-008-044-MY3) (Y.T.H). This work is also supported by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division under contract no. DE-AC02-76SF00515 (synthesis and transport measurements) as well as SuperC and the Kavli Foundation (in situ transport) (Y.L., Y.T., Y.Y., H.Y.H.), and by Grants-in-Aid for Scientific Research (KAKENHI, grant numbers 22H00104 and 25H00600) from the Japanese Society for the Promotion of Science (JSPS) (Y.K., V.S.). Part of this work is supported by the US Department of Energy (DOE), Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division, under contract DE-AC02-76SF00515 (B.Y.W., Z.X.S). We further acknowledge support of the HLD at HZDR, member of the European Magnetic Field Laboratory (EMFL) as well as support under the European Union’s Horizon 2020 research and innovation programme through the ISABEL project (No. 871106).